Calculate the pH of the resulting solution if 31.0 mL of 0.310 M HCI(aq) is added to (a) 41.0 mL of 0.310 M NaOH(aq). Number (b) 21.0 mL of 0.410 M NaOH(aq). Number
Answer #1

HCl + NaOH --------> NaCl + H2O

millimoles of HCL = 31.0 x 0.310 = 9.61

a) millimoles of NaOH added = 41.0 x 0.310 = 12.71

12.71 - 9.61 = 3.1 millimoles NaOH left

[NaOH] = 3.1 / 31+41 = 0.043 M

[OH-] = [NaOH] = 0.043 M

pOH = - log [OH-]

pOH = - log [0.043]

pOH = 1.37

pH = 14 - 1.37

pH = 12.63

b) millimoles of NaOH = 21 x0 .410 = 8.61

9.61 - 8.61 = 1.0 millimoles HCl left

[HCl] = 1.0 / 31 + 21 = 0.019 M

[H+] = [HCl] = 0.019 M

pH = - log [H+]

pH = - log [0.019]

pH = 1.72